OpenStudy (anonymous):

what is the solution of n^2-49=0

OpenStudy (anonymous):

squareroot of n^2 =49

OpenStudy (anonymous):

square root of 49 is 7

OpenStudy (anonymous):

no not the squareroot of 49 the square root of a number squared (n^2) equals 49

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

the solution is n=49

OpenStudy (anonymous):

because you are solving you would add 49 to both sides because you are trying to get n alone and you would get n^2=49 and then you would need to find the squareroot of n to get n alone so you would get n=49

OpenStudy (anonymous):

wrote the equation wrong. its ... n^2 -49=0

OpenStudy (anonymous):

is it -7, 7, \[\pm7\], or no solution?

OpenStudy (anonymous):

i didn't mention the equation acutally